OpenStudy (anonymous):

simplify the following expression -220x^3y^2/44x^10y^9

OpenStudy (anonymous):

first deal with terms in x: -220x^3 / 44 x^10 = -5 x^-7 terms in y : y^2 / y^ 9 = y ^ -7 = -5 x^-7 / y^-7 = = - y^7 / 5 x^7

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol the one i just done or anothter

OpenStudy (anonymous):

hmm its not one of my choices

OpenStudy (anonymous):

ok - i might have made a mistake what are the choices?

OpenStudy (anonymous):

1) 5/x^7^7

OpenStudy (anonymous):

2) -5/x^7y^7

OpenStudy (anonymous):

3) 220/x^7y^7 4)-5x^13y^11

OpenStudy (anonymous):

did u get that?

OpenStudy (anonymous):

yes i'm just checking them out

OpenStudy (anonymous):

oh ok soorry

OpenStudy (anonymous):

yes sorry guy - i just done it on paper - its awkward on the screen - i did go wrong - the answer is -5 / x^7y^7 Number 1 do u need me to write it out in detail?

OpenStudy (anonymous):

-220x^3 / 44 x^10 = -5/ x^7 y^2 / y^ 9 =1/y^7 multiplying: = -5 /x^7y^7

OpenStudy (anonymous):

no -its number 2!! -5

OpenStudy (anonymous):

lol ok:)

OpenStudy (anonymous):

got there in the end - maybe i should turn in!!